Tormo bought a perfectly square table but found it wont fit through his door. He reduces its width by 9cm but also extends its length by 9cm- its new area is 309cm^2. what was the area of his original square table.
Answer by Boreal(15235)

original table has side length x and area x^2 cm^2.
(x-9)(x+9)=309 cm^2
x^2-81=309
x^2=390, and the area of the original table was 390 cm^2. ANSWER
x=19.75 cm and the new table was 10.75 cm x 28.75 cm